# Unsupervised Representation Learning from Sparse Transformation Analysis

### v2

@@ Appendix C Proof of how the Hamilton-Jacobi equation solves Optimal Transport
Theorem C.1 (Benamou-Brenier Formula [benamou2000computational]).
@@ D.1 Model Architectures
For the MLPs that parameterize the scalar potential $u({\bm{z}},t)$ and the divergence-free vector field ${\bm{r}}({\bm{z}})$, we use linear layers to encode the latent samples and use Tanh as the activation function. The sinusoidal positional embeddings [vaswani2017attention] are used to embed the timestep $t$. For our variational auto-encoders, the encoder simply takes four stacked convolution layers with ReLU activation functions, while the decoder consists of four transposed convolution layers.
@@ D.3 Training Details
On MNIST [lecun1998mnist], the training process lasts $50,000$ iterations. We only train the spike component in the first $20,000$ iterations and then integrate the slab component into the training for the rest of the iterations. On Shapes3D [3dshapes18], we set the total training iterations to $100,000$ and split the iterations fifty-fifty into the two training stages. The batch size is set to $128$ and $64$ on MNIST and Shapes3D, respectively. We use the Adam optimizer with a learning rate of 1e41e{-}4. The input images are of the size $28{\times}28$ on MNIST and of the size $64{\times}64$ on Shapes3D. The quantitative results are reported as mean $\pm$ standard deviation computer over $5$ runs with random initialization.
@@ E.1 Disentanglement Metrics
As discussed in [Tzelepis_2021_ICCV, song2023flow], following $\beta$-VAE [higgins2016beta], there are many disentanglement metrics proposed for single-dimension traversal methods, such as DCI [eastwood2018framework] and MIG [chen2018isolating]. These metrics assume that each latent dimension corresponds to one generative factor and manipulating these single dimensions would trigger distinct output transformations. The recent disentanglement methods [shen2021closed, Tzelepis_2021_ICCV, song2023householder] propose a more realistic disengagement setting: all the latent dimensions are perturbed by vector arithmetic for output variations. If one evaluates these vector-based disentanglement methods using the metrics designed for single-dimension manipulations, their scores would drop considerably and are not comparable. Nonetheless, certain disentanglement metrics such as the VP score [zhu2020learning] can be adopted for the evaluation of general disentanglement baselines as they do not pose any assumptions on the latent space. Instead, the VP metric takes a lightweight neural network to learn classifying a dataset of different image pairs ${\bm{x}}_{0},{\bm{x}}_{T}]$ under the few-shot learning setting (i.e., only $1\%$ or $10\%$ of the dataset is used as the training set). The validation accuracy reflects the distinguishability of these learned traversal directions, which is a reasonable surrogate for the disentanglement score.
@@ E.2 Impact of Divergence-free Vector Fields
TABLE XII: Approximation error of each divergence-free vector field on MNIST [lecun1998mnist].
@@ E.3 Traversal for More Steps
Fig. 15 displays the traversals of our learned flow fields for more time points (longer sequences). We see that our STA still allows for smooth interpolation across timesteps. Notice that the traversal step is even larger than the maximal sequence length of Falcol3D and Issac3D [nie2020semi], but the latent flows keep smooth transitions during time evolution.

### v3

@@ Appendix C Proof of how the Hamilton-Jacobi equation solves Optimal Transport
Theorem C.1 (Benamou-Brenier Formula [7]).
@@ D.1 Model Architectures
For the MLPs that parameterize the scalar potential $u({\bm{z}},t)$ and the divergence-free vector field ${\bm{r}}({\bm{z}})$, we use linear layers to encode the latent samples and use Tanh as the activation function. The sinusoidal positional embeddings [91] are used to embed the timestep $t$. For our variational auto-encoders, the encoder simply takes four stacked convolution layers with ReLU activation functions, while the decoder consists of four transposed convolution layers.
@@ D.3 Training Details
On MNIST [56], the training process lasts $50,000$ iterations. We only train the spike component in the first $20,000$ iterations and then integrate the slab component into the training for the rest of the iterations. On Shapes3D [11], we set the total training iterations to $100,000$ and split the iterations fifty-fifty into the two training stages. The batch size is set to $128$ and $64$ on MNIST and Shapes3D, respectively. We use the Adam optimizer with a learning rate of 1e41e{-}4. The input images are of the size $28{\times}28$ on MNIST and of the size $64{\times}64$ on Shapes3D. The quantitative results are reported as mean $\pm$ standard deviation computer over $5$ runs with random initialization.
@@ E.1 Disentanglement Metrics
As discussed in [90, 79], following $\beta$-VAE [38], there are many disentanglement metrics proposed for single-dimension traversal methods, such as DCI [27] and MIG [13]. These metrics assume that each latent dimension corresponds to one generative factor and manipulating these single dimensions would trigger distinct output transformations. The recent disentanglement methods [75, 90, 82] propose a more realistic disengagement setting: all the latent dimensions are perturbed by vector arithmetic for output variations. If one evaluates these vector-based disentanglement methods using the metrics designed for single-dimension manipulations, their scores would drop considerably and are not comparable. Nonetheless, certain disentanglement metrics such as the VP score [106] can be adopted for the evaluation of general disentanglement baselines as they do not pose any assumptions on the latent space. Instead, the VP metric takes a lightweight neural network to learn classifying a dataset of different image pairs ${\bm{x}}_{0},{\bm{x}}_{T}]$ under the few-shot learning setting (i.e., only $1\%$ or $10\%$ of the dataset is used as the training set). The validation accuracy reflects the distinguishability of these learned traversal directions, which is a reasonable surrogate for the disentanglement score.
@@ E.2 Impact of Divergence-free Vector Fields
TABLE XII: Approximation error of each divergence-free vector field on MNIST [56].
@@ E.3 Traversal for More Steps
Fig. 15 displays the traversals of our learned flow fields for more time points (longer sequences). We see that our STA still allows for smooth interpolation across timesteps. Notice that the traversal step is even larger than the maximal sequence length of Falcol3D and Issac3D [62], but the latent flows keep smooth transitions during time evolution.